OpenStudy (anonymous):

I am working with point slope formula and am not sure how I set it up my points are 2,3, and 6,7, with a slope of one.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

i know that but what is x and y. How do i get the intercepts.

OpenStudy (anonymous):

y1=3, m=1, x1=2

OpenStudy (anonymous):

solve for y. y=x+1

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

so to get intercepts, set one or other =0. so for yint, set x=0.

OpenStudy (anonymous):

so y would be 1.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for x in, set y=0

OpenStudy (anonymous):

so for that half of the equation it would be 1-2?

OpenStudy (anonymous):

so x=-1

OpenStudy (anonymous):

why does x equal -1

OpenStudy (anonymous):

ok your equation is y=x+1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i guess.

OpenStudy (anonymous):

what is the y part of the equation. i don't get it @floritttt

OpenStudy (anonymous):

when u have y= to something

OpenStudy (anonymous):

that is your equation for the line

OpenStudy (anonymous):

would my equation be (1-3)=1(-1-2)

OpenStudy (anonymous):

how.

OpenStudy (anonymous):

@floritttt